THEORETICAL PREDICTIONS OF THE STRUCTURES AND ENERGETICS OF ClF_{n}^{+/−}(n=1−5) IONS: EXTENDED STUDIES OF HYPERVALENT SPECIES USING THE RECOUPLED PAIR BONDING MODEL

\underline{L. CHEN}, D. E. WOON, and T. H. DUNNING, JR., Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

Following studies of the nature of covalent and recoupled pair bonding in PF_{n}(n=1−5), SF_{n}(n=1−6), and ClF_{n}(n=1−7), we have used the recoupled pair bonding model to investigate the ClF_{n}^{+/−}(n=1−5) ions. The behavior of the cation series parallels the SF_{n} series to great extent, while the behavior of the anion series would parallel that of ArF_{n} if those species were bound. In the present research, we predict experimentally measurable properties, such as the electron affinities (EA) and ionization energies (IE) of the associated neutral ClF_{n} species. The optimized ground state structures, bond energies, and spectral properties of ClF_{n}^{+/−} were obtained by employing high level ab initio calculations (MRCI, CCSD(T)) with correlation consistent basis sets. Our predictions agree well with the structures of experimentally observed species, including ClF_{2}^{+}, ClF_{2}^{−}, ClF_{4}^{+}, and ClF_{4}^{−}. Low-lying excited states of ClF^{+}(^{4}Σ^{−}) and ClF_{2}^{+}(^{3}B_{1}), were also identified. We found the excited ^{3}B_{1} state of ClF_{2}^{+} is bent with an angle of 154.8^{∘} and a bond length of 1.596 Å at the RCCSD(T)/AVQZ level. We also found a major difference between SF_{2} and ClF_{2}^{+}: the ^{3}A_{2} state of SF_{2} is bound while the same state of ClF_{2}^{+} is repulsive. Second, we systematically explored the bond formation processes, adding F atoms one at a time to the optimized ClF_{n}^{+/−}(1≤n≤5) ions. We found the ionization energies and the electronic affinities both exhibit an oscillating trend as seen in the bond energies of the neutral ClF_{n} species. The structural and energetic data obtained by our calculations might help spectroscopists to identify new ClF_{n}^{+/−} species or to detect new electronic states of known species.